OpenStudy (anonymous):

What is the slope of a line perpendicular to a line that contains the points (-8, -4) and (0, -8)?

OpenStudy (***[isuru]***):

first find the slope of the line where u got the coordinates already.. \[m = \frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\]\[m = \frac{ -4 +8 }{ -8 } =- \frac{ 1}{ 2 }\]

OpenStudy (***[isuru]***):

now... if the slope of a line is "m" then the slope of the line which is perpendicular to is \[-\frac{ 1 }{ m }\]

OpenStudy (***[isuru]***):

now if we consider \[m = -\frac{ 1 }{ 2 }\] then the slope of the line we want is \[-\frac{ 1 }{ m } = - (-\frac{ 2 }{ 1 }) = +2\] got it :) ?